Solve the given integral equation or integro-differential equation for y(t). y (1)+ Ja-vy(v) dv=5t, y(0) = 0 0 y(t)=D
Answer #1

410204 hiven qw+ !* (tv) yu dv : 17 both sich Apply Laplace transform tyl1t} + 14 d * 14 - 6) giu) dv} 7463 Sicht- 400) + 1 YCompany - temo O=A+B BE-=-113 B = -1/3 compainz constats termo I=A+C (=1-A c=1-1 3 c=213 from a Q V3 -15 + 73 = + 33+1 3+1 sالا) 3 t 나 일 (5)+(2) 일회분 ㅗㅗ 부 -일) + 3 t 3 (S) +1 both sides laplace trong tom Apply inverse = 음행해로 일일 얼 발행 et sin J3 r eft ca